Please help me how to solve this.
The product of two consecutive positive even integers is 120. Find the integers.

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Please help me how to solve this.
The product of two consecutive positive even integers is 120. Find the integers.
-----------
Hint: sqrt(120) =~ 11

Answer by ikleyn(52798)

.
Amswer. 10 and 12.

Solution

x*(x+2) = 120,

x^2 + 2x - 120 = 0.

Factor left side

(x-10)*(x+12) = 0.

The only solution in the positive domain is x=10, so the answer is 10 and the numbers are 10 and 12.